OpenStudy (anonymous):

Iss the following always, sometimes or never true? X + 1 = X + 0

OpenStudy (anonymous):

always not true

OpenStudy (anonymous):

1 does not equal to zero

OpenStudy (anonymous):

:) Thankssss. I got it.

OpenStudy (mathteacher1729):

You can think of this in terms of money. If I have x dollars, and you have x + 1 dollars, how could we ever have the same amount of money? (If I have 5 dollars, and you have 5+ 1 = 6 dollars, or if I have 100 dollars and you have 100+1 = 101 dollars, etc.) You will always have one dollar more. Algebraically we start with x + 0 = x + 1 ignore the plus zero x = x + 1 subtract x from both sides 0 = 1 ... um. Zero does not equal one. :) So the statement is NEVER true.

OpenStudy (anonymous):

Thank you, I never understood these problems because it's just numbers and letters pulled together. However I do understand now. It's easier once it's explained in money terms. :)

OpenStudy (mathteacher1729):

Yay, glad I could help! :D